OpenStudy (anonymous):

Evaluate each expression. a. log₂32 b. log(.0001) c. lne⁻³ d. log₄1

OpenStudy (anonymous):

You have to use the properties, for example, in a, your 32 is a power of 2, in b the same, 0,0001 is a power of 10, ln=log_e and in d, there's only one way to obtain 1 in a number that is not 1 xD (a^0=1)

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

okay, if you have, for example log 100 it is 2, right?, because 10^2=100, that's because of logarithmic function definition when you have \[a^x=b\] that's equal to \[\log_a{b}=x \] so, in a you have \[\log_2 {32}=\log_2 {2^5}\] that means that x=5, got it?

OpenStudy (anonymous):

oh okay got that

OpenStudy (anonymous):

i just dont understand how to find the final answer

OpenStudy (anonymous):

the "x" in ths case is the final answer

OpenStudy (anonymous):

so for a. the answer is x=5?

OpenStudy (anonymous):

the answer is 5, i used the x to compare it with the definition

OpenStudy (anonymous):

oh okay great, thanks!! how about b?

OpenStudy (anonymous):

i got -4 for b

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

goood

OpenStudy (anonymous):

i didn't get c

OpenStudy (anonymous):

or d :( those two confuse me

OpenStudy (anonymous):

c is the easier one \[\ln b=\log_e{b} \] that means \[\ln e^{-3}= \log_e {e^{-3}}\]

OpenStudy (anonymous):

so is logee−3 the final answer?

OpenStudy (anonymous):

the way you typed it of course

OpenStudy (anonymous):

nope, that means that is equivalent, now if \[\log_2 {2^5} =5\] then \[\log_e {e^{-3}}=-3\] got that?

OpenStudy (anonymous):

ohh okay so then it would be -3?

OpenStudy (anonymous):

ohh okay i sort of understand now

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

how about the last one

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

4^0=1, then your answer is?

OpenStudy (anonymous):

1? @Umangiasd

OpenStudy (anonymous):

nope, look at what you have in the question.